OpenStudy (truwhovian11):

sos will MEDAL and FAN Which percent is equivalent to 5 1/20 ? A. 5.05% B. 5.12% C. 505% D. 520%

OpenStudy (truwhovian11):

@wilsondanielle @Tal_Tal @ayeedomo

OpenStudy (anonymous):

i tell you

OpenStudy (truwhovian11):

OK thats fine

OpenStudy (wilsondanielle):

start by finding the decimal - and converting to an improper fraction. (5*20)+1/20 = 101/20 = 5.05 5.05*100 = 505%

OpenStudy (truwhovian11):

ok is the improper fraction is 101/2o

OpenStudy (anonymous):

OK you WON

OpenStudy (wilsondanielle):

The decimal is over 1 and that means the percent is over 100%. because you have more than 20/20 (aka 1)

OpenStudy (truwhovian11):

what do u mean i won

OpenStudy (wilsondanielle):

but the overall process (multiplying/dividing by 100%) is still the same.

OpenStudy (anonymous):

to wilsondaniel cause he gave you the answer

OpenStudy (truwhovian11):

um still confused sorry its 6:00 pm and i need sleep

OpenStudy (ayeedomo):

Do you still need help with this?

OpenStudy (anonymous):

LOL

OpenStudy (truwhovian11):

yes

OpenStudy (anonymous):

yes he does

OpenStudy (truwhovian11):

im a she thank you very much

OpenStudy (wilsondanielle):

I'm a she... My name is Danielle hence my username. Not Daniel

OpenStudy (truwhovian11):

i know thx

OpenStudy (wilsondanielle):

The improper fraction is 101/20. Find the decimal (either through cross multiplication like you saw on the last question or through a calculator) and then convert by multiplying by 100%.

OpenStudy (ayeedomo):

So you have 5 1/20 Since 5 is a whole number, you'll have a 5 before the decimal. One easy way to calculate the decimal is to make the fraction out of 100. 20 goes into 100 5 times, so multiply by 5. 1/20 = 5/100 you can read 5/100 as "five hundredths". So the answer is "five and five hundredths" or 5.05

OpenStudy (truwhovian11):

oh thx um since @ayeedomo is the only one here ill give you the meadal

OpenStudy (truwhovian11):

can u help with one more

OpenStudy (ayeedomo):

Sure XD

OpenStudy (truwhovian11):

ill make another question